# Bruhat operads

### v1

@@ 1 Introduction
The aim of the present paper is to introduce some planar operads build from higher Bruhat orders, cf. [2].
Recall that a planar111sometimes called nonsymmetric operad of sets $\mathcal{O}$ consists of a family of sets $\{\mathcal{O}(n)\}_{n\in\mathbb{Z},n\geq 1}$, together with composition maps
@@
and a unit $1\in\mathcal{O}(1)$ satisfying the associativity and the unit axiom as for a usual operad, except that we do not require an action of symmetric groups $S(n)$ on $\mathcal{O}(n)$, cf. [5], Def. 3.12 and Def. 2.1 below.
@@
Let $d,n$ be integers, $1\leq d\leq n$. In [2] (cf. also [3]) certain posets $B(n,d)$ called higher Bruhat orders have been introduced, their definition is recalled below in 5.1.
One can describe them roughly as follows. $B(n,1)=S(n)$ with the classical weak Bruhat order. The poset $S(n)$ admits unique maximal and minimal elements, namely, the minimal element is the unit permutation $e$, whereas the maximal one is the permutation $\sigma_{max}$ of maximal length222where length means the usual length $\ell$ in a Coxeter group, so $\ell(\sigma_{max})=n/(n-1)/2$.. By definition $B(n,1)$ is certain quotient of the set of maximal (i.e. noncompressible) chains in $S(n)$ connecting $e$ with $\sigma_{max}$.
Afterwards one proceeds by induction on $d$. Each $B(n,d)$ is a poset admitting unique minimal and maximal elements, say $b_{min}^{d}$ and $b_{max}^{d}$, and $B(n,d+1)$ is certain quotient of the set of maximal chains in $B(n,d)$ connecting $b_{min}^{d}$ with $b_{max}^{d}$.
A remarkable paper by Ziegler [6] contains a number of deep constructions and results on higher Bruhat orders. They are essential in our study.
@@ 4 Symmetric groups
For any set $[n+\sum_{i=0}^{n}k_{i}]$ we consider permutations $\sigma\in S(n+\sum_{i}k_{i})$ which permutes only teams and are identical on the idles.
@@ 4.3 Weak Bruhat order
Both operads $S$ and $F$ are compatible with Bruhat orders. Namely
@@ 5 Higher Bruhat orders and insertions
Let us recall some definitions from [2].
@@
The set $B(n,d)$ of Bruhat orders for (${[n]}\choose d$ is defined as the quotient $A(n,d)/\sim$ where $\sim$ is an equivalence relation described in [2], no 2.
Each $B(n,d)$ is a poset, to be denoted $\mathcal{HB}^{d}(n)$. One of main results of [2] states that any order from $\mathcal{HB}^{d}(n)$ can be included into a maximal chain joining the minimal order with the maximal one.
@@ 5.5 Ziegler criterion
Ziegler [6], Th. 4.1(B), characterised the subsets of (${[n]}\choose{d+1}$ which appear as sets of inversions of Bruhat orders.
@@ 5.6.1 Example
(the union is disjoint). Then we say that

### v2

@@ 1 Introduction
The aim of the present paper is to introduce some planar operads build from higher Bruhat orders, cf. [MSFA].
Recall that a planar (sometimes called nonsymmetric) operad of sets $\mathcal{O}$ consists of a family of sets $\{\mathcal{O}(n)\}_{n\in\mathbb{Z},n\geq 1}$, together with composition maps
@@
and a unit $1\in\mathcal{O}(1)$ satisfying the associativity and the unit axiom as for a usual operad, except that we do not require an action of symmetric groups $S(n)$ on $\mathcal{O}(n)$, cf. [M], Def. 3.12 and Def. 2.1 below.
@@
Let $d,n$ be integers, $1\leq d\leq n$. In [MSFA] (cf. also [MS]) certain posets $B(n,d)$ called higher Bruhat orders have been introduced, their definition is recalled below in 5.1.
One can describe them roughly as follows. $B(n,1)=S(n)$ with the classical weak Bruhat order. The poset $S(n)$ admits unique maximal and minimal elements, namely, the minimal element is the unit permutation $e$, whereas the maximal one is the permutation $\sigma_{max}$ of maximal length111where length means the usual length $\ell$ in a Coxeter group, so $\ell(\sigma_{max})=n/(n-1)/2$.. By definition $B(n,2)$ is a certain quotient of the set of maximal (i.e. noncompressible222which means a collection $e<x_{1}<\ldots<x_{n}<\sigma_{max}$ which cannot be included into a longer collection) chains in $S(n)$ connecting $e$ with $\sigma_{max}$.
Afterwards one proceeds by induction on $d$. Each $B(n,d)$ is a poset admitting unique minimal and maximal elements, say $b_{min}^{d}$ and $b_{max}^{d}$, and $B(n,d+1)$ is a certain quotient of the set of maximal chains in $B(n,d)$ connecting $b_{min}^{d}$ with $b_{max}^{d}$.
A remarkable paper by Ziegler [Z] contains a number of deep constructions and results on higher Bruhat orders. They are essential in our study.
@@ 4 Symmetric groups
For any set $[n+\sum_{i=0}^{n}k_{i}]$ we consider permutations $\sigma\in S(n+\sum_{i}k_{i})$ which permute only nuclei and are identical on the electrons.
@@ 4.3 Weak Bruhat order
For example the weak Bruhat order for $S(3)$ looks as follows:
$e<(12)<(123)<(13)$ and $e<(23)<(132)<(13)$.
Both operads $S$ and $F$ are compatible with weak Bruhat orders. Namely
@@ 5 Higher Bruhat orders and insertions
Let us recall some definitions from [MSFA].
@@
The set $B(n,d)$ of Bruhat orders for (${[n]}\choose d$ is defined as the quotient $A(n,d)/\sim$ where $\sim$ is an equivalence relation described in [MSFA], no 2.
Each $B(n,d)$ is a poset, to be denoted $\mathcal{HB}^{d}(n)$. One of main results of [MSFA] states that any order from $\mathcal{HB}^{d}(n)$ can be included into a maximal chain joining the minimal order with the maximal one.
@@ 5.5 Ziegler criterion
Ziegler [Z], Th. 4.1(B), characterised the subsets of (${[n]}\choose{d+1}$ which appear as sets of inversions of Bruhat orders.
@@ 5.6 Main construction: insertions of higher Bruhat orders
(the union is disjoint).